# Two-dimensional delta Bose gas in a weighted space

### v1

@@ Acknowledgment
The comparison argument in Section 2 applies here, too. Consider the multiplicative operator ~$\widetilde{\mathcal{E}}:L^{2}_{a}(\mathbb{R}^{4})\to L^{2}_{a}(\mathbb{R}^{4})$, (~$(\widetilde{\mathcal{E}}\psi)(y_{\mathrm{r}},y_{\mathrm{c}}):=e^{a\sum_{\sigma=\pm 1}|y_{\mathrm{c}}+\frac{\varepsilon\sigma}{2}y_{\mathrm{r}}|_{1}}\psi(y_{\mathrm{r}},y_{\mathrm{c}})$. The bounds in (2.14) continue to hold when we replace $\mathcal{P}_{\alpha}(t)$, $\mathcal{P}_{\alpha}(t)^{*}$, $\mathcal{P}_{\alpha\alpha^{\prime}}(t)$, $\mathcal{J}_{\alpha}(t)$, $\mathcal{E}^{2}$, and $\mathcal{E}^{-2}$ with $\mathcal{P}^{\varepsilon}_{\alpha}(t)$, $\mathcal{P}^{\varepsilon}_{\alpha}(t)^{*}$, $\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)$, $\mathcal{J}^{\varepsilon}_{\alpha}(t)$, $\widetilde{\mathcal{E}}$, and ~$\widetilde{\mathcal{E}}^{-1}$, respectively. Next, the same comparison argument leading to (2.16) applies also to their $\varepsilon$ analogs and gives, for $\alpha\neq\alpha^{\prime}$,
$\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(2t)\big{\|}_{2\to 2}\ ,$ (3.9a)
$\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(t)^{*}\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(2t)^{*}\big{\|}_{2\to 2}\ ,$ (3.9b)
$\big{\|}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(2t)\big{\|}_{2\to 2}\ ,$ (3.9c)
$\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-(c+2)t}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)\Big{\|}_{2,a\to 2,a}$ $\leq c\,\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-2t}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(2t)\Big{\|}_{2\to 2}\ .$ (3.9d)
$\big{\|}\mathcal{J}^{\varepsilon}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\,\|\mathsf{j}^{\varepsilon}(t)\|_{2,a\to 2,a}\,\|\mathsf{p}(t)\|_{2\to 2}\cdot\|\mathsf{p}(2t)\|_{2\to 2}^{n-2}\ .$ (3.9e)
@@
Further, since $\phi$ has a compact support, $\|\mathsf{j}^{\varepsilon}(t)\|_{2,a\to 2,a}\leq c\|\mathsf{j}^{\varepsilon}(t)\|_{2\to 2}$. Using this bound and $\|\mathsf{p}(s)\|_{2\to 2}\leq 1$ turns (3.9e) into $\|\mathcal{J}^{\varepsilon}_{\alpha}(t)\|_{2,a\to 2,a}\leq ce^{ct}\,\|\mathsf{j}^{\varepsilon}(t)\|_{2,a\to 2,a}$. Given these comparison bounds, the main task is to prove the following bounds.
$\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(t)\big{\|}_{2\to 2}=\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(t)^{*}\big{\|}_{2\to 2}$ $\leq c\,t^{-1/2}\ ,$ (3.10a)
$\big{\|}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)\big{\|}_{2\to 2}$ $\leq c\,t^{-1}\ ,$ (3.10b)
$\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-t}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)\Big{\|}_{2\to 2}$ $\leq c\ ,$ (3.10c)
$\big{\|}\mathsf{j}^{\varepsilon}(t)\big{\|}_{2\to 2}$ $\leq c\,t^{-1}\big{|}\log\big{(}t\wedge\tfrac{1}{2}\big{)}\big{|}^{-2}\cdot e^{c\,t}\ .$ (3.10d)
@@
Postponing the proof of (3.10) to Section 3.1, we now complete the proof of Theorem 1.1(2). First, combining (3.10) and Lemma 2.1 for $c_{1}=\beta_{\varepsilon}$ and for a large enough $\lambda$ gives the first statement in Theorem 1.1(2). To prove the second statement, let $(\mathbf{1}_{\leq M}f)(x):=f(x)\mathbf{1}_{|x|_{1}\leq M}$, let $\mathbf{1}_{>M}:=\mathbf{1}-\mathbf{1}_{\leq M}$, and write
$\begin{split}\big{\|}&(\mathcal{Q}(t)-\mathcal{Q}^{\varepsilon}(t))f\big{\|}_{2,a}\leq\big{\|}\mathbf{1}_{\leq M}(\mathcal{Q}(t)-\mathcal{Q}^{\varepsilon}(t))\mathbf{1}_{\leq M^{\prime}}f\big{\|}_{2,a}\\ &+\sum_{\mathcal{T}=\mathcal{Q}(t),\mathcal{Q}^{\varepsilon}(t)}\big{\|}\mathbf{1}_{>M}\mathcal{T}\mathbf{1}_{\leq M^{\prime}}f\big{\|}_{2,a}+\sum_{\mathcal{T}=\mathcal{Q}(t),\mathcal{Q}^{\varepsilon}(t)}\big{\|}\mathcal{T}\mathbf{1}_{>M^{\prime}}f\big{\|}_{2,a}\ .\end{split}$ (3.11)
@@
3.1. Proving (3.10)
The bounds (3.10a)–(3.10c) follow from existing results. For $\alpha=ij$, let $N_{\alpha}$ denote the map $\mathbb{R}^{2[\![n]\!]}\to\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha},N_{\alpha}p:=(p_{i}+p_{j},p_{[\![n]\!]\setminus\alpha})$, and let $p^{-}_{\alpha}:=(p_{i}-p_{j})/2$. For $f\in L^{2}(\mathbb{R}^{2[\![n]\!]})$ and $g\in L^{2}(\mathbb{R}^{4}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha})$, let
$\widehat{f}(p):=\int_{\mathbb{R}^{2[\![n]\!]}}\frac{\mathrm{d}p}{(2\pi)^{n}}e^{-\mathbf{i}p\cdot x}f(x)\ ,$ $\overbracket{g}(y_{\mathrm{r}},q):=\prod_{i\in\mathrm{c}\cup[\![n]\!]\setminus\alpha}\int_{\mathbb{R}^{2}}\frac{\mathrm{d}y_{i}}{2\pi}e^{-\mathbf{i}q_{i}\cdot y_{i}}\cdot g(y)$ (3.12)
@@
$\big{\langle}g,\mathcal{P}^{\varepsilon}_{\alpha}(t)f\big{\rangle}$ $=\int_{\mathbb{R}^{2}}\frac{\mathrm{d}y_{\mathrm{r}}}{2\pi}\int_{\mathbb{R}^{2[\![n]\!]}}\mathrm{d}p\,\overline{\overbracket{g}(y_{\mathrm{r}},N_{\alpha}p)}\,\phi(y_{\mathrm{r}})e^{\mathbf{i}\varepsilon y_{\mathrm{r}}\cdot p^{-}_{\alpha}}\,e^{-\frac{t}{2}|p|_{2}^{2}}\widehat{f}(p)\ ,$ (3.13)
@@
$\begin{split}\big{\langle}g_{1},\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)g_{2}\big{\rangle}&=\int_{\mathbb{R}^{2}}\frac{\mathrm{d}y_{\mathrm{r}}}{2\pi}\int_{\mathbb{R}^{2}}\frac{\mathrm{d}y_{\mathrm{r}}^{\prime}}{2\pi}\int_{\mathbb{R}^{2[\![n]\!]}}\mathrm{d}p\\ &\cdot\overline{\overbracket{g_{1}}(y_{\mathrm{r}},N_{\alpha}p)}\,\phi(y_{\mathrm{r}})\,e^{-\frac{t}{2}|p|_{2}^{2}+\mathbf{i}\varepsilon(y_{\mathrm{r}}\cdot p^{-}_{\alpha}-y_{\mathrm{r}}^{\prime}\cdot p^{-}_{\alpha^{\prime}})}\phi(y_{\mathrm{r}}^{\prime})\,\overbracket{g_{2}}(y_{\mathrm{r}}^{\prime},N_{\alpha^{\prime}}p)\ .\end{split}$ (3.14)
@@
In (3.13), bound $|\phi(y_{\mathrm{r}})e^{\mathbf{i}\varepsilon y_{\mathrm{r}}\cdot p^{-}_{\alpha}}|\leq c\,\phi(y_{\mathrm{r}})$ and $e^{-t|p|_{2}^{2}/2}\leq e^{-t|p_{i}-p_{j}|_{2}^{2}/4}$, write the resulting integrand as the product of $c\,\phi(y_{\mathrm{r}})\,\overline{\overbracket{g}(y_{\mathrm{r}},N_{\alpha}p)}e^{-t|p_{i}-p_{j}|_{2}^{2}/4}$ and $\widehat{f}(p)$, and apply the Cauchy–Schwarz inequality over the $p$ integral. Doing so gives the bound $c\,t^{-1/2}\,(\int_{\mathbb{R}^{2}}\mathrm{d}y_{\mathrm{r}}\phi(y_{\mathrm{r}})\|g(y_{\mathrm{r}},\cdot)\|_{2})\cdot\|f\|_{2}$. Further applying the Cauchy–Schwarz inequality over the $y_{\mathrm{r}}$ integral gives (3.10a). Applying a similar argument to (3.14) gives (3.10b). Applying $\int_{0}^{\infty}\mathrm{d}t\,e^{-t}$ to (3.14) followed by using [15, Equations (3.1), (3.3), (3.4), (3.6)] gives (3.10c).
@@
Proving (3.10d) requires more work. We begin with a preliminary bound. Let $\|\cdot\|_{\mathrm{HS}}$ denote the Hilbert–Schmidt norm and let $b(u):=\int_{\mathbb{R}^{4}}\mathrm{d}y_{\mathrm{r}}\mathrm{d}y_{\mathrm{r}}^{\prime}\,\phi^{2}(y_{\mathrm{r}})e^{-u|y_{\mathrm{r}}-y_{\mathrm{r}}^{\prime}|_{2}^{2}/4}\phi^{2}(y_{\mathrm{r}}^{\prime})$. We have $\|\phi\mathsf{p}_{\varepsilon}(2\tau)\phi\|_{2\to 2}\leq\|\phi\mathsf{p}_{\varepsilon}(2\tau)\phi\|_{\mathrm{HS}}=b(\varepsilon^{2}/\tau)^{1/2}/4\pi\tau$. Recall that $\phi$ is $C^{\infty}$ smooth, has a compact support, and satisfies $\int_{\mathbb{R}^{2}}\mathrm{d}x\,\phi^{2}(x)=1$. Using these properties, it is not hard to check that $b$ is $C^{\infty}$ smooth and strictly decreasing on $u\in[0,\infty)$, that $b(0)=1$, that $b^{\prime}(0)<0$, and that $\limsup_{u\to\infty}b(u)u<\infty$. These properties together imply that $b(u)\leq(1+c_{0}u)^{-1}$, for some $c_{0}\in(0,\infty)$ depending only on $\phi$. Hence $\|\phi\mathsf{p}_{\varepsilon}(2\tau)\phi\|_{2\to 2}\leq(4\pi)^{-1}(\tau^{2}+c_{0}\varepsilon^{2}\tau)^{-1/2}$. Using this in (3.4) gives
@@
$\|\mathsf{j}^{\varepsilon}(t)\|_{2,a\to 2,a}\leq B_{\varepsilon}(t)\ ,\qquad B_{\varepsilon}(t):=\beta_{\varepsilon}\sum_{\ell=1}^{\infty}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{\ell}\frac{\beta_{\varepsilon}}{4\pi\sqrt{\tau_{k}^{2}+c_{0}\varepsilon^{2}\tau_{k}}}\ .$ (3.15)
@@
The bound (3.15) reduces our task to bounding $B_{\varepsilon}(t)$. Let us consider the cases $t\leq\varepsilon^{2}c_{0}$ and $t>\varepsilon^{2}c_{0}$ separately. For $t\leq\varepsilon^{2}c_{0}$, bound $1/\sqrt{\tau_{k}^{2}+c_{0}\varepsilon^{2}\tau_{k}}\leq 1/\sqrt{\varepsilon^{2}\tau_{k}}$ and use the Dirichlet integral formula $\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{\ell}\tau_{k}^{-1/2}=t^{\ell/2-1}\Gamma(1/2)^{\ell}/\Gamma(\ell/2)$ to get
$B_{\varepsilon}(t)\leq c\,\beta_{\varepsilon}\,t^{-1}\,\,\sum_{\ell=1}^{\infty}\frac{(c\beta_{\varepsilon}\sqrt{t\varepsilon^{-2}})^{\ell}}{\Gamma(\ell/2)}=c\,\beta_{\varepsilon}\,t^{-1/2}\varepsilon^{-1}\,\,\sum_{\ell=1}^{\infty}\frac{(c\beta_{\varepsilon})^{\ell}(\sqrt{t\varepsilon^{-2}})^{\ell-1}}{\Gamma(\ell/2)}\ .$ (3.16)
@@
The series converges for all small enough $\varepsilon$ because $\beta_{\varepsilon}\to 0$. Using $t\varepsilon^{-2}\leq c_{0}$ and $\beta_{\varepsilon}\leq c/|\log\varepsilon|$, we bound $B_{\varepsilon}(t)$ by $c\,t^{-1/2}\cdot\varepsilon^{-1}|\log\varepsilon|^{-2}$. Since $t^{-1/2}(\log t)^{-2}$ decreases in $t$ for small enough $t$ and since $t\leq\varepsilon^{2}c_{0}$, the last expression is bounded by $c\,t^{-1/2}\cdot t^{-1/2}|\log t|^{-2}$.
@@
Moving on to the case $t>\varepsilon^{2}c_{0}$, we begin by deriving a contour-integral formula of $B_{\varepsilon}(t)$. Consider the Laplace transform of the factor in the product in (3.15):
$C_{\varepsilon}(\lambda):=\int_{0}^{\infty}\mathrm{d}s\,\frac{e^{-\lambda s}\beta_{\varepsilon}}{4\pi\sqrt{s^{2}+c_{0}\varepsilon^{2}s}}\ .$ (3.17)

### v2

@@ 1 Introduction
The comparison argument in Section 2 applies here, too. Consider the multiplicative operator ~$\widetilde{\mathcal{E}}:L^{2}_{a}(\mathbb{R}^{4})\to L^{2}_{a}(\mathbb{R}^{4})$, (~$(\widetilde{\mathcal{E}}\psi)(y_{\mathrm{r}},y_{\mathrm{c}}):=e^{a\sum_{\sigma=\pm 1}|y_{\mathrm{c}}+\frac{\varepsilon\sigma}{2}y_{\mathrm{r}}|_{1}}\psi(y_{\mathrm{r}},y_{\mathrm{c}})$. The bounds in (18) continue to hold when we replace $\mathcal{P}_{\alpha}(t)$, $\mathcal{P}_{\alpha}(t)^{*}$, $\mathcal{P}_{\alpha\alpha^{\prime}}(t)$, $\mathcal{J}_{\alpha}(t)$, $\mathcal{E}^{2}$, and $\mathcal{E}^{-2}$ with $\mathcal{P}^{\varepsilon}_{\alpha}(t)$, $\mathcal{P}^{\varepsilon}_{\alpha}(t)^{*}$, $\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)$, $\mathcal{J}^{\varepsilon}_{\alpha}(t)$, $\widetilde{\mathcal{E}}$, and ~$\widetilde{\mathcal{E}}^{-1}$, respectively. Next, the same comparison argument leading to (20) applies also to their $\varepsilon$ analogs and gives, for $\alpha\neq\alpha^{\prime}$,
$\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(2t)\big{\|}_{2\to 2}\ ,$ (29a)
$\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(t)^{*}\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(2t)^{*}\big{\|}_{2\to 2}\ ,$ (29b)
$\big{\|}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(2t)\big{\|}_{2\to 2}\ ,$ (29c)
$\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-(c+2)t}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)\Big{\|}_{2,a\to 2,a}$ $\leq c\,\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-2t}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(2t)\Big{\|}_{2\to 2}\ .$ (29d)
$\big{\|}\mathcal{J}^{\varepsilon}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\,\|\mathsf{j}^{\varepsilon}(t)\|_{2,a\to 2,a}\,\|\mathsf{p}(t)\|_{2\to 2}\cdot\|\mathsf{p}(2t)\|_{2\to 2}^{n-2}\ .$ (29e)
@@
Further, since $\phi$ has a compact support, $\|\mathsf{j}^{\varepsilon}(t)\|_{2,a\to 2,a}\leq c\|\mathsf{j}^{\varepsilon}(t)\|_{2\to 2}$. Using this bound and $\|\mathsf{p}(s)\|_{2\to 2}\leq 1$ turns (29e) into $\|\mathcal{J}^{\varepsilon}_{\alpha}(t)\|_{2,a\to 2,a}\leq ce^{ct}\,\|\mathsf{j}^{\varepsilon}(t)\|_{2,a\to 2,a}$. Given these comparison bounds, the main task is to prove the following bounds.
$\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(t)\big{\|}_{2\to 2}=\big{\|}\mathcal{P}^{\varepsilon}_{\alpha}(t)^{*}\big{\|}_{2\to 2}$ $\leq c\,t^{-1/2}\ ,$ (30a)
$\big{\|}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)\big{\|}_{2\to 2}$ $\leq c\,t^{-1}\ ,$ (30b)
$\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-t}\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)\Big{\|}_{2\to 2}$ $\leq c\ ,$ (30c)
$\big{\|}\mathsf{j}^{\varepsilon}(t)\big{\|}_{2\to 2}$ $\leq c\,t^{-1}\big{|}\log\big{(}t\wedge\tfrac{1}{2}\big{)}\big{|}^{-2}\cdot e^{c\,t}\ .$ (30d)
@@
Postponing the proof of (30) to Section 3.1, we now complete the proof of Theorem 1.1(2). First, combining (30) and Lemma 2.1 for $c_{1}=\beta_{\varepsilon}$ and for a large enough $\lambda$ gives the first statement in Theorem 1.1(2). To prove the second statement, let $(\mathbf{1}_{\leq M}f)(x):=f(x)\mathbf{1}_{|x|_{1}\leq M}$, let $\mathbf{1}_{>M}:=\mathbf{1}-\mathbf{1}_{\leq M}$, and write
$\begin{split}\big{\|}&(\mathcal{Q}(t)-\mathcal{Q}^{\varepsilon}(t))f\big{\|}_{2,a}\leq\big{\|}\mathbf{1}_{\leq M}(\mathcal{Q}(t)-\mathcal{Q}^{\varepsilon}(t))\mathbf{1}_{\leq M^{\prime}}f\big{\|}_{2,a}\\ &+\sum_{\mathcal{T}=\mathcal{Q}(t),\mathcal{Q}^{\varepsilon}(t)}\big{\|}\mathbf{1}_{>M}\mathcal{T}\mathbf{1}_{\leq M^{\prime}}f\big{\|}_{2,a}+\sum_{\mathcal{T}=\mathcal{Q}(t),\mathcal{Q}^{\varepsilon}(t)}\big{\|}\mathcal{T}\mathbf{1}_{>M^{\prime}}f\big{\|}_{2,a}\ .\end{split}$ (31)
@@
3.1 Proving (30)
The bounds (30a)–(30c) follow from existing results. For $\alpha=ij$, let $N_{\alpha}$ denote the map $\mathbb{R}^{2[\![n]\!]}\to\mathbb{R}^{2}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha},N_{\alpha}p:=(p_{i}+p_{j},p_{[\![n]\!]\setminus\alpha})$, and let $p^{-}_{\alpha}:=(p_{i}-p_{j})/2$. For $f\in L^{2}(\mathbb{R}^{2[\![n]\!]})$ and $g\in L^{2}(\mathbb{R}^{4}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha})$, let
$\widehat{f}(p):=\int_{\mathbb{R}^{2[\![n]\!]}}\frac{\mathrm{d}p}{(2\pi)^{n}}e^{-\mathbf{i}p\cdot x}f(x)\ ,$ $\overbracket{g}(y_{\mathrm{r}},q):=\prod_{i\in\mathrm{c}\cup[\![n]\!]\setminus\alpha}\int_{\mathbb{R}^{2}}\frac{\mathrm{d}y_{i}}{2\pi}e^{-\mathbf{i}q_{i}\cdot y_{i}}\cdot g(y)$ (32)
@@ 3.1 Proving (30)
$\big{\langle}g,\mathcal{P}^{\varepsilon}_{\alpha}(t)f\big{\rangle}$ $=\int_{\mathbb{R}^{2}}\frac{\mathrm{d}y_{\mathrm{r}}}{2\pi}\int_{\mathbb{R}^{2[\![n]\!]}}\mathrm{d}p\,\overline{\overbracket{g}(y_{\mathrm{r}},N_{\alpha}p)}\,\phi(y_{\mathrm{r}})e^{\mathbf{i}\varepsilon y_{\mathrm{r}}\cdot p^{-}_{\alpha}}\,e^{-\frac{t}{2}|p|_{2}^{2}}\widehat{f}(p)\ ,$ (33)
@@
$\begin{split}\big{\langle}g_{1},\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t)g_{2}\big{\rangle}&=\int_{\mathbb{R}^{2}}\frac{\mathrm{d}y_{\mathrm{r}}}{2\pi}\int_{\mathbb{R}^{2}}\frac{\mathrm{d}y_{\mathrm{r}}^{\prime}}{2\pi}\int_{\mathbb{R}^{2[\![n]\!]}}\mathrm{d}p\\ &\cdot\overline{\overbracket{g_{1}}(y_{\mathrm{r}},N_{\alpha}p)}\,\phi(y_{\mathrm{r}})\,e^{-\frac{t}{2}|p|_{2}^{2}+\mathbf{i}\varepsilon(y_{\mathrm{r}}\cdot p^{-}_{\alpha}-y_{\mathrm{r}}^{\prime}\cdot p^{-}_{\alpha^{\prime}})}\phi(y_{\mathrm{r}}^{\prime})\,\overbracket{g_{2}}(y_{\mathrm{r}}^{\prime},N_{\alpha^{\prime}}p)\ .\end{split}$ (34)
@@
In (33), bound $|\phi(y_{\mathrm{r}})e^{\mathbf{i}\varepsilon y_{\mathrm{r}}\cdot p^{-}_{\alpha}}|\leq c\,\phi(y_{\mathrm{r}})$ and $e^{-t|p|_{2}^{2}/2}\leq e^{-t|p_{i}-p_{j}|_{2}^{2}/4}$, write the resulting integrand as the product of $c\,\phi(y_{\mathrm{r}})\,\overline{\overbracket{g}(y_{\mathrm{r}},N_{\alpha}p)}e^{-t|p_{i}-p_{j}|_{2}^{2}/4}$ and $\widehat{f}(p)$, and apply the Cauchy–Schwarz inequality over the $p$ integral. Doing so gives the bound $c\,t^{-1/2}\,(\int_{\mathbb{R}^{2}}\mathrm{d}y_{\mathrm{r}}\phi(y_{\mathrm{r}})\|g(y_{\mathrm{r}},\cdot)\|_{2})\cdot\|f\|_{2}$. Further applying the Cauchy–Schwarz inequality over the $y_{\mathrm{r}}$ integral gives (30a). Applying a similar argument to (34) gives (30b). Applying $\int_{0}^{\infty}\mathrm{d}t\,e^{-t}$ to (34) followed by using [15, Equations (3.1), (3.3), (3.4), (3.6)] gives (30c).
@@
Proving (30d) requires more work. We begin with a preliminary bound. Let $\|\cdot\|_{\mathrm{HS}}$ denote the Hilbert–Schmidt norm and let $b(u):=\int_{\mathbb{R}^{4}}\mathrm{d}y_{\mathrm{r}}\mathrm{d}y_{\mathrm{r}}^{\prime}\,\phi^{2}(y_{\mathrm{r}})e^{-u|y_{\mathrm{r}}-y_{\mathrm{r}}^{\prime}|_{2}^{2}/4}\phi^{2}(y_{\mathrm{r}}^{\prime})$. We have $\|\phi\mathsf{p}_{\varepsilon}(2\tau)\phi\|_{2\to 2}\leq\|\phi\mathsf{p}_{\varepsilon}(2\tau)\phi\|_{\mathrm{HS}}=b(\varepsilon^{2}/\tau)^{1/2}/4\pi\tau$. Recall that $\phi$ is $C^{\infty}$ smooth, has a compact support, and satisfies $\int_{\mathbb{R}^{2}}\mathrm{d}x\,\phi^{2}(x)=1$. Using these properties, it is not hard to check that $b$ is $C^{\infty}$ smooth and strictly decreasing on $u\in[0,\infty)$, that $b(0)=1$, that $b^{\prime}(0)<0$, and that $\limsup_{u\to\infty}b(u)u<\infty$. These properties together imply that $b(u)\leq(1+c_{0}u)^{-1}$, for some $c_{0}\in(0,\infty)$ depending only on $\phi$. Hence $\|\phi\mathsf{p}_{\varepsilon}(2\tau)\phi\|_{2\to 2}\leq(4\pi)^{-1}(\tau^{2}+c_{0}\varepsilon^{2}\tau)^{-1/2}$. Using this in (24) gives
@@
$\|\mathsf{j}^{\varepsilon}(t)\|_{2,a\to 2,a}\leq B_{\varepsilon}(t)\ ,\qquad B_{\varepsilon}(t):=\beta_{\varepsilon}\sum_{\ell=1}^{\infty}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{\ell}\frac{\beta_{\varepsilon}}{4\pi\sqrt{\tau_{k}^{2}+c_{0}\varepsilon^{2}\tau_{k}}}\ .$ (35)
@@
The bound (35) reduces our task to bounding $B_{\varepsilon}(t)$. Let us consider the cases $t\leq\varepsilon^{2}c_{0}$ and $t>\varepsilon^{2}c_{0}$ separately. For $t\leq\varepsilon^{2}c_{0}$, bound $1/\sqrt{\tau_{k}^{2}+c_{0}\varepsilon^{2}\tau_{k}}\leq 1/\sqrt{\varepsilon^{2}\tau_{k}}$ and use the Dirichlet integral formula $\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{\ell}\tau_{k}^{-1/2}=t^{\ell/2-1}\Gamma(1/2)^{\ell}/\Gamma(\ell/2)$ to get
$B_{\varepsilon}(t)\leq c\,\beta_{\varepsilon}\,t^{-1}\,\,\sum_{\ell=1}^{\infty}\frac{(c\beta_{\varepsilon}\sqrt{t\varepsilon^{-2}})^{\ell}}{\Gamma(\ell/2)}=c\,\beta_{\varepsilon}\,t^{-1/2}\varepsilon^{-1}\,\,\sum_{\ell=1}^{\infty}\frac{(c\beta_{\varepsilon})^{\ell}(\sqrt{t\varepsilon^{-2}})^{\ell-1}}{\Gamma(\ell/2)}\ .$ (36)
@@
The series converges for all small enough $\varepsilon$ because $\beta_{\varepsilon}\to 0$. Using $t\varepsilon^{-2}\leq c_{0}$ and $\beta_{\varepsilon}\leq c/|\log\varepsilon|$, we bound $B_{\varepsilon}(t)$ by $c\,t^{-1/2}\cdot\varepsilon^{-1}|\log\varepsilon|^{-2}$. Since $t^{-1/2}(\log t)^{-2}$ is decreasing for $t\leq\varepsilon^{2}c_{0}$ as $\varepsilon\to 0$, the last expression is bounded by $c\,t^{-1/2}\cdot t^{-1/2}|\log t|^{-2}$.
@@
Moving on to the case $t>\varepsilon^{2}c_{0}$, we begin by deriving a contour-integral formula of $B_{\varepsilon}(t)$. Consider the Laplace transform of the factor in the product in (35):
$C_{\varepsilon}(\lambda):=\int_{0}^{\infty}\mathrm{d}s\,\frac{e^{-\lambda s}\beta_{\varepsilon}}{4\pi\sqrt{s^{2}+c_{0}\varepsilon^{2}s}}\ .$ (37)